OpenStudy (anonymous):

"Determine all positive n integers such that n divides by 3^n-2^n." Halp! I don't know where to start. :c

OpenStudy (campbell_st):

looks like mathematical Induction Step 1 prove true for n = 1 3^1 -2^1 = 1 and n =1 is a factor of 1. Step 2: assume true for n= k then 3^k - 2^k = km where m is a factor or m = (3^k - 2^k)/k Step 3: Prove for n=k+1 then 3^(k+1) - 2^(k+1) =3 x 3^k - 2x2^k ummm.. a bit stuck

OpenStudy (asnaseer):

Assuming the question is to find all positive integer solutions to:\[p=\frac{n}{3^n-2^n}\]then we know one trivial solution (as pointed out by campbell above) is \(n=1\). we can show that this is the ONLY solution by proving (using induction) that:\[3^n-2^n>n\hspace{2cm}\text{for all }n\ge2\]firstly we can show this is true for \(n=2\):\[3^2-2^2=9-4=5\gt2\]next we assume it is true for \(n=k\):\[3^k-2^k\gt k\implies3^k\gt k+2^k\tag{a}\]and now we just need to prove it is also true for \(n=k+1\):\[\begin{align} 3^{k+1}-2^{k+1}&=3\cdot3^k-2\cdot2^k\\ &\gt3(k+2^k)-2\cdot2^k\hspace{2cm}\text{using (a)}\\ &\gt3k+3\cdot2^k-2\cdot2^k\\ &\gt3k+2^k\\ &\gt k+(2k+2^k)\\ &\gt k+1 \end{align}\] Hence the only solution is \(n=1\).

OpenStudy (anonymous):

Clever. :o

OpenStudy (asnaseer):

thx